Jefferson Lab HALL A Logbook for December 99 - All EXCEPT Automatic Entries

  • Main INDEX

    Thursday

  • 31926 12/23/99 10:27 Saha Beam Parameters at End of Run
  • 31924 12/23/99 07:29 M. Rvachev XX century shift summary
  • 31923 12/23/99 07:12 x. jiang number of events in kinematic 26
  • 31922 12/23/99 07:02 M. Rvachev Beam off
  • 31918 12/23/99 06:30 M. Rvachev Reboot hiresxt20
  • 31914 12/23/99 05:45 x. jiang beam position plot since 22;50 last night.
  • 31913 12/23/99 05:42 x. jiang magnet window at run 1600.
  • Wednesday

  • 31879 12/22/99 23:35 LeRose Beam Strip Charts
  • 31876 12/22/99 23:16 LeRose OTR
  • 31875 12/22/99 23:09 LeRose OTR Filter
  • 31864 12/22/99 20:55 Nilanga analysis of #26 data (cont.)
  • 31863 12/22/99 20:49 Nilanga Analysis of #26 data
  • 31861 12/22/99 20:35 LeRose OTR
  • 31860 12/22/99 20:34 LeRose Beam Diagnostics
  • 31859 12/22/99 20:29 LeRose Beam Strip tool
  • 31854 12/22/99 19:11 LeRose Problem with 1574
  • 31847 12/22/99 18:29 LeRose Magnets etc.
  • 31841 12/22/99 17:25 LeRose Beam Back
  • 31840 12/22/99 17:22 LeRose Injector Problems
  • 31839 12/22/99 17:09 LeRose Beam Strip Tool
  • 31835 12/22/99 16:48 LeRose More Strip Tools
  • 31834 12/22/99 16:45 LeRose Strip Tool
  • 31830 12/22/99 16:35 LeRose Magnets Ready
  • 31829 12/22/99 16:25 LeRose Quad cycling
  • 31828 12/22/99 16:04 S. Gilad morning shift 12/22
  • 31826 12/22/99 15:48 Templon new trigger downloaded
  • 31825 12/22/99 15:37 Templon changing hadron angle
  • 31820 12/22/99 14:57 Templon prescale factors nilanga pointing
  • 31812 12/22/99 14:31 Templon run sheet problem
  • 31806 12/22/99 13:40 Templon inelastic link point short run
  • 31803 12/22/99 13:25 R. Michaels deadtime copying to MSS
  • 31801 12/22/99 12:56 Templon nilanga pointing 12C analysis
  • 31800 12/22/99 12:35 templon OTR logging briefly fubared
  • 31796 12/22/99 12:13 Templon, Liy kin 16 analysis
  • 31795 12/22/99 12:01 Higinbotham Kin#27 Missing Energy
  • 31793 12/22/99 11:40 Eugene Beam optics stripchart
  • 31792 12/22/99 11:36 Eugene Magnet stripchart
  • 31789 12/22/99 11:16 Templon data analysis for kin 16
  • 31785 12/22/99 10:19 Saha Arc-Integral measurement
  • 31778 12/22/99 09:03 Saha EZ_DPP screen for Kinematics #16
  • 31777 12/22/99 09:01 Saha HAC_xt screen for Kinematics #16
  • 31776 12/22/99 09:00 Saha e89044_beam summary
  • 31773 12/22/99 08:34 Saha Beam parameters summary
  • 31768 12/22/99 08:05 M. Rvachev Shift summary
  • 31763 12/22/99 07:44 M. Rvachev setspec and spectrometer angle
  • 31762 12/22/99 07:18 M. Rvachev Beam current durung shift
  • 31761 12/22/99 07:15 fatiha elastic rate
  • 31757 12/22/99 06:45 M. Rvachev Magnet settings
  • 31746 12/22/99 05:23 M. Rvachev High deadtime
  • 31735 12/22/99 04:15 x. jiang elastic rate at 4.044 GeV
  • Tuesday

  • 31714 12/21/99 22:46 Higinbotham Accelerator Still Down
  • 31713 12/21/99 21:50 Higinbotham Strip Charts
  • 31708 12/21/99 18:07 Higinbotham Accelerator Trouble
  • 31707 12/21/99 17:27 Higinbotham New Settings
  • 31706 12/21/99 17:22 Higinbotham Electron Spectrometer
  • 31705 12/21/99 17:02 suleiman Beam Summary
  • 31697 12/21/99 16:13 suleiman magnet summary
  • 31695 12/21/99 16:09 S. Gilad No beam until 15:45
  • 31692 12/21/99 15:48 Nilanga 3He elastic rate
  • 31691 12/21/99 15:46 Templon beam is back, nilanga pointing
  • 31687 12/21/99 12:39 Templon cryotarg pc rebooted
  • 31686 12/21/99 11:11 R. Michaels RE: SPOT vs MCC beam size
  • 31685 12/21/99 10:48 R. Michaels advice about hac
  • 31684 12/21/99 10:08 Templon why we want absolute locks
  • 31683 12/21/99 09:40 templon test entry from hac/adaql1
  • 31682 12/21/99 09:25 Saha Beam parameters
  • 31681 12/21/99 09:24 Saha Arc-Integral measurement, Dec 20, 1999
  • 31679 12/21/99 08:21 Z. Chai Beam energy stripplot
  • 31678 12/21/99 08:14 M. Rvachev Current stripchart
  • 31677 12/21/99 08:11 M. Rvachev Shift summary
  • 31673 12/21/99 07:48 M. Rvachev All feedbacks off
  • 31671 12/21/99 07:23 M. Rvachev runs 1502-1513 in ep tune (junk?)
  • 31669 12/21/99 07:09 M. Rvachev absolute energy fast feedback off
  • 31664 12/21/99 06:50 M. Rvachev Raster size
  • 31663 12/21/99 06:48 M. Rvachev End of run comments changed
  • 31655 12/21/99 06:06 Z. Chai E-Arm D1 N2 low level
  • 31652 12/21/99 05:52 Z. Chai Magnet Setting
  • 31651 12/21/99 05:41 A. Deur eP DAQ trouble.
  • 31639 12/21/99 02:01 M. Rvachev EQ3 quench
  • 31633 12/21/99 00:23 Higinbotham New Spectrometer Settings
  • Monday

  • 31631 12/20/99 23:49 Higinbotham 4 GeV Run Plan
  • 31627 12/20/99 22:41 Eugene HAC cleanup
  • 31626 12/20/99 21:44 S. Gilad Summary
  • 31625 12/20/99 21:03 Higinbotham 4 GeV BDL Check
  • 31618 12/20/99 17:34 Templon cryotarg alarm setpt change
  • 31615 12/20/99 17:30 Templon target change
  • 31612 12/20/99 17:09 templon roc 14 crash, bad EOR
  • 31611 12/20/99 16:21 Saha Arc-Integral measurement
  • 31610 12/20/99 16:15 Templon cryotarg temp excursion
  • 31607 12/20/99 14:48 S. Gilad DAQ crash
  • 31605 12/20/99 13:59 S. Gilad beam parameters
  • 31599 12/20/99 11:21 S. Gilad Beam optics
  • 31598 12/20/99 11:14 S. Gilad beam parameters
  • 31597 12/20/99 11:10 S. Gilad Magnet stripchart
  • 31591 12/20/99 08:50 S. Gilad Beam parameters
  • 31590 12/20/99 08:25 G. M. Urciuo owl shift beam optic stripcharts
  • 31588 12/20/99 07:51 G.M.Urciuoli 112-20-99 owl shift summary.
  • 31585 12/20/99 07:28 voutier Run 1480 statistics
  • 31584 12/20/99 06:58 G. M. Urciuo beam optic stripcharts - part 2.
  • 31583 12/20/99 06:56 G. M. Urciuo beam optic stripcharts - part 1.
  • 31580 12/20/99 06:40 R. Michaels Trigger Pattern (high stats)
  • 31579 12/20/99 06:26 R. Michaels synch checks
  • 31578 12/20/99 05:59 R. Michaels Trigger Latch Pattern
  • 31576 12/20/99 05:49 R. Michaels BPM snapshots
  • 31574 12/20/99 04:55 voutier Run 1477 statistics
  • 31571 12/20/99 04:29 voutier Target Chart
  • 31570 12/20/99 04:08 R. Michaels Why T2 and T4 are background
  • 31569 12/20/99 03:37 R. Michaels Trigger Stability
  • 31566 12/20/99 03:10 R. Michaels New detmap (S0 added)
  • 31565 12/20/99 03:04 voutier Run 1476 statistics
  • 31564 12/20/99 02:48 R. Michaels S0 Data for T3, T4
  • 31561 12/20/99 01:47 voutier Run 1475 satistics
  • 31558 12/20/99 01:12 G. M. Urciuo owl shift beam optic stripcharts part 2
  • 31557 12/20/99 01:05 G. M Urciuol 12-20-99: owl shift beam optic stripcharts
  • 31556 12/20/99 00:50 voutier Run 1474 statistics
  • 31555 12/20/99 00:26 Suleiman Beam parameters
  • Sunday

  • 31553 12/19/99 23:52 suleiman halog
  • 31551 12/19/99 23:38 Suleiman shift Summary
  • 31546 12/19/99 21:33 fatiha coincidence events Run 1470
  • 31539 12/19/99 19:18 fatiha cryotarg status
  • 31536 12/19/99 17:59 Higinbotham Shift Summary
  • 31532 12/19/99 16:43 Templon please check e-arm vdc ps
  • 31531 12/19/99 16:20 templon cryotarg stripcharts
  • 31526 12/19/99 15:55 Templon E arm vdc low voltages
  • 31527 12/19/99 15:55 Templon ROC again
  • 31521 12/19/99 14:41 R. Michaels meaning of (-1,-1) in T2
  • 31518 12/19/99 14:03 Templon trigger check
  • 31516 12/19/99 13:35 Templon roc problems solved
  • 31513 12/19/99 13:08 Templon Hall Access b/c ROC problems
  • 31512 12/19/99 12:40 Templon daq problems (dead ROCs)
  • 31507 12/19/99 10:43 Templon cryotarget shiftly stripcharting
  • 31504 12/19/99 09:19 Templon trigger checks ooops
  • 31503 12/19/99 09:18 Templon VDC double tracks
  • 31502 12/19/99 09:10 Templon trigger checks
  • 31496 12/19/99 08:02 G. Quemener owl shift summary
  • 31492 12/19/99 07:20 E. Cisbani Kin 27 e_miss and p_miss spectra
  • 31491 12/19/99 06:51 R. Michaels Stability of Trigger
  • 31488 12/19/99 06:26 R. Michaels Snapshot of BPMs each event
  • 31484 12/19/99 04:27 R. Michaels scaler charge overflows
  • 31483 12/19/99 04:10 E. Cisbani Kin 27 (e,e'p) Coincidence Rate approx. 270/h
  • 31480 12/19/99 03:10 quemener beam strip charts
  • 31479 12/19/99 03:09 quemener beam strip charts
  • 31478 12/19/99 02:57 adaq Magnet summary window
  • 31475 12/19/99 01:53 R. Michaels detector checks
  • 31472 12/19/99 00:05 F. GARIBALDI Shift summary
  • Saturday

  • 31468 12/18/99 23:44 F. Garibaldi run 1943
  • 31463 12/18/99 22:02 Garibaldi Optics strip chart
  • 31462 12/18/99 21:59 Garibaldi OTR
  • 31461 12/18/99 21:50 F. Garibaldi run 1442
  • 31456 12/18/99 18:38 adaq rscaler to check rates
  • 31455 12/18/99 18:34 garibaldi OTR strip chart
  • 31452 12/18/99 17:58 Higinbotham Kin 27 Approx( 36 e'p / hr )
  • 31447 12/18/99 16:31 Templon daq problems continue
  • 31444 12/18/99 16:05 Saha Arc BdL measurement
  • 31437 12/18/99 15:41 Templon 3He(e,e'p) test run finished
  • 31435 12/18/99 15:35 Higinbotham (e,e'p) Event Cuts
  • 31432 12/18/99 15:22 Saha e89044-beam stripchart
  • 31431 12/18/99 15:15 Templon electrons are back
  • 31428 12/18/99 15:01 Templon prescale factor change
  • 31425 12/18/99 14:58 Templon daq anomalies (?)
  • 31412 12/18/99 13:03 Higinbotham CrossNormalization File1423
  • 31411 12/18/99 12:47 Templon carbon inelastic nilanga pointing
  • 31409 12/18/99 12:44 Templon new trigger settings downloaded, kin 27
  • 31408 12/18/99 12:43 Templon trigger check hana
  • 31406 12/18/99 12:28 Higinbotham Kin 27 Spect. Settings
  • 31405 12/18/99 12:15 Templon target change He->carbon
  • 31403 12/18/99 12:10 Templon adaqs2 session crash
  • 31400 12/18/99 11:48 epstein, tem config change, kin 27 production
  • 31398 12/18/99 10:59 Templon trigger check kin 27 link point
  • 31394 12/18/99 10:37 Templon cryotarget, IOC reboot, temp weirdness
  • 31393 12/18/99 10:29 Higinbotham E-HRS to 87.5 and 0.583 Mev
  • 31391 12/18/99 09:33 Epstein, tem e spect change, new espace hdr
  • 31387 12/18/99 08:07 R. Michaels How to Download Trigger til Christmas
  • 31386 12/18/99 08:03 R. Gilman owl shift summary
  • 31384 12/18/99 08:00 R. Michaels Kin 17 started
  • 31381 12/18/99 07:28 R. Michaels BPM 12 (again)
  • 31376 12/18/99 07:14 R. Michaels BPM 8 and 12 into ROC13
  • 31370 12/18/99 04:48 R. Michaels Six hours is about right
  • 31369 12/18/99 04:36 R. Michaels fraction of anomolous T2,T4 (not too bad)
  • 31366 12/18/99 04:20 garibaldi optics chart
  • 31365 12/18/99 04:17 garibaldi optics strip chart
  • 31364 12/18/99 04:15 F. Garibaldi magnet stripchart
  • 31363 12/18/99 04:14 GaRIBAldi otr screen
  • 31358 12/18/99 02:06 R. Michaels rates on he-3 elastic
  • 31357 12/18/99 01:50 R. Michaels he-3 elastic spectra
  • 31356 12/18/99 01:43 gilman B dl measurement
  • 31351 12/18/99 00:37 Quemener swing shift summary
  • Friday

  • 31348 12/17/99 23:26 Adev spot output for Q2=19.39
  • 31341 12/17/99 22:54 voutier New Field Settings
  • 31333 12/17/99 19:34 quemener strip charts 2
  • 31332 12/17/99 19:33 quemener beam strip chart 1
  • 31330 12/17/99 18:24 Epstein Simulation of kin. #17
  • 31324 12/17/99 17:08 K. Aniol reproduciblitiy of Q2=5.01 /fm2
  • 31319 12/17/99 16:02 Z. Chai Beam Current Strip plot
  • 31318 12/17/99 16:00 Z. Chai Shift Summary
  • 31313 12/17/99 11:53 Higinbotham Rate at 15.0 fm^-2
  • 31312 12/17/99 11:48 Z. Chai Beam Current, Beam Position
  • 31311 12/17/99 11:33 Z. Chai Magnets setting
  • 31310 12/17/99 11:29 Z. Chai Beam Current Stripplot
  • 31309 12/17/99 11:26 Doug Rate at 5.01 fm^-2
  • 31308 12/17/99 11:03 M Jones add spectrometer pointing variable to data stream
  • 31307 12/17/99 10:48 Doug Rate at 15.0 fm^-2
  • 31305 12/17/99 08:49 R. Michaels T2 -- seem to be background
  • 31301 12/17/99 08:29 M. Rvachev Shift summary
  • 31297 12/17/99 07:38 E. Cisbani energy stripchart
  • 31296 12/17/99 07:32 R. Iommi magnet screenshot (corrected) at Q2=15.00
  • 31295 12/17/99 07:31 E. Cisbani Elastic 3He, Q2=15.0
  • 31294 12/17/99 07:29 E. Cisbani Elastic 3He, Q2=5.01
  • 31293 12/17/99 07:27 E. Cisbani Elastic 3He, Q2=1.31
  • 31292 12/17/99 07:25 E. Cisbani Elastic 3He Q2=1.9, Plots
  • 31291 12/17/99 07:20 E. Cisbani Stripcharts
  • 31290 12/17/99 07:18 E. Cisbani Beamline stripcharts
  • 31287 12/17/99 06:25 R.Iommi magnet screenshot
  • 31278 12/17/99 05:22 R. Iommi Magnet screenshot
  • 31277 12/17/99 05:16 M. Rvachev Beam position screen
  • 31269 12/17/99 04:31 R.Iommi Magnet screen
  • 31258 12/17/99 03:39 R. Iommi Magnet screen
  • 31251 12/17/99 02:46 R. Michaels T4 problem -- partly resolved
  • Thursday

  • 31235 12/16/99 21:38 K. Aniol revised elastic normalization point
  • 31234 12/16/99 17:27 R. Michaels T4 from run 1349.
  • 31232 12/16/99 16:36 R. Michaels T2,T4 rates. Check of hardware
  • 31229 12/16/99 15:05 Epstein Revised kinematics 17,27,16
  • 31228 12/16/99 12:16 R. Michaels problem of T4 - can affect normalization
  • 31226 12/16/99 07:39 kuss target loop 1 fan 6%
  • 31225 12/16/99 06:46 Suleiman Helium Elastic Calibration Data
  • 31223 12/16/99 06:31 suleiman 3He Charge Form Factor
  • 31210 12/16/99 04:41 M. Rvachev Beam position
  • 31188 12/16/99 01:16 Epstein Emiss for kin. #17
  • 31187 12/16/99 01:08 R. Iommi Beeper
  • 31183 12/16/99 00:14 kuss target stripcharts
  • 31179 12/16/99 00:00 R. Roche Shift Summary 16:00-24:00
  • Wednesday

  • 31175 12/15/99 23:35 R Roche Spot++ for run #1346
  • 31165 12/15/99 22:41 R Roche Harp scan
  • 31161 12/15/99 21:42 M. Rvachev Analisys problem
  • 31160 12/15/99 21:19 R Roche Magnet Strip Charts
  • 31159 12/15/99 20:57 Saha Arc-Energy at 4 GeV, Dec 15,1999
  • 31158 12/15/99 20:49 Epstein Quad cycling
  • 31157 12/15/99 20:46 perdrisat going back to 842
  • 31156 12/15/99 20:25 perdrisat CODA turned on.
  • 31154 12/15/99 19:21 R Roche Beam optics strip charts #2
  • 31153 12/15/99 19:20 R Roche Beam optics strip charts #1
  • 31152 12/15/99 19:00 K. Aniol correction to th_e for kin 17'
  • 31151 12/15/99 18:48 K. Aniol kin #17' set up
  • 31149 12/15/99 16:15 Saha Broken wire for Scanner #6
  • 31148 12/15/99 16:10 Saha Strip charts, BPMs and Current, Dec 15, 0800-1600
  • 31147 12/15/99 16:09 Saha Strip charts, Energy, Dec 15 0800-1600
  • 31146 12/15/99 16:08 Saha Shift summary 08:00- 16:00
  • 31145 12/15/99 14:12 R. Michaels re: DAQ problems
  • 31136 12/15/99 13:21 kuss target stripcharts
  • 31128 12/15/99 10:41 Saha Problem wirth Scanner # 6
  • 31126 12/15/99 09:15 Margaziotis Harp scans
  • 31124 12/15/99 08:56 Margaziotis Dummy for I-scan
  • 31116 12/15/99 08:15 G. M. Urciuo 12-15-99: owl shift report.
  • 31106 12/15/99 07:21 G.M.Urciuoli Spot for run 1323
  • 31101 12/15/99 06:49 M. Rvachev Spot for run 1321
  • 31094 12/15/99 06:09 G. M. Urciuo Spot for 1318
  • 31087 12/15/99 05:33 M. Rvachev Spot for run 1312
  • 31081 12/15/99 05:06 G. Urciuoli OTR
  • 31079 12/15/99 04:54 G. M. Urciuo Spot for 1312
  • 31073 12/15/99 04:20 M. Rvachev Spot for 1310
  • 31070 12/15/99 04:13 M. Rvachev Spot for 1309
  • 31063 12/15/99 03:50 Adev Spot size
  • 31055 12/15/99 03:33 E. Cisbani Raster Oscilloscope Inputs
  • 31046 12/15/99 01:02 kuss target stripcharts
  • 31045 12/15/99 00:47 Eric Voutier Binders
  • 31044 12/15/99 00:47 kuss Hall access to replace a faulty BCM power supply
  • Tuesday

  • 31040 12/14/99 23:39 kuss target heating study
  • 31039 12/14/99 23:33 R. Roche magnet strip charts
  • 31038 12/14/99 23:32 R. Roche Magnet Strip Charts
  • 31036 12/14/99 23:08 voutier Harp Scans
  • 31035 12/14/99 22:54 kuss BCM trips
  • 31034 12/14/99 21:54 Epstein run plan for 12/14/99
  • 31030 12/14/99 21:12 M. Rvachev Start of run comments
  • 31029 12/14/99 19:56 Eric Voutier New field settings
  • 31028 12/14/99 19:56 D. Margaziot Status
  • 31027 12/14/99 18:02 D. Margaziot System check
  • 31024 12/14/99 17:44 D. Margaziot Run 1293 - junk
  • 31022 12/14/99 17:25 K. Aniol current elastic settings for 3He
  • 31021 12/14/99 08:17 R. Michaels Rate versus current - summary
  • 31020 12/14/99 08:14 K. Aniol beam use
  • 31017 12/14/99 07:49 R. Michaels temp vs time at 50 microamps
  • 31008 12/14/99 06:46 R. Michaels beam returns, table of results
  • 30998 12/14/99 06:22 R. Michaels more detailed elastic spectra
  • 30997 12/14/99 05:50 R. Michaels lead target, 5 microamps
  • 30996 12/14/99 05:36 R. Michaels beam trips at 30 micro
  • 30990 12/14/99 05:14 R. Michaels Elastic Peak seen
  • 30982 12/14/99 04:48 R. Michaels EPICS screen, lead test
  • 30980 12/14/99 04:44 R. Michaels beam spot
  • 30974 12/14/99 04:20 R. Michaels Lead Target test
  • 30973 12/14/99 04:04 R. Michaels T4 pattern seems wrong
  • 30969 12/14/99 03:13 K. Aniol end 4He, start Pb tgt tests
  • 30964 12/14/99 02:06 K. Aniol setting low momenta in hrse
  • 30963 12/14/99 01:44 K. Aniol start run 1270
  • 30961 12/14/99 01:29 K. Aniol CODA end and Deadtime
  • Monday

  • 30951 12/13/99 23:01 epstein I scan
  • 30945 12/13/99 22:22 epstein hadron arm momentum change
  • 30939 12/13/99 21:56 kuss strip chart
  • 30937 12/13/99 21:55 kuss strip chart
  • 30934 12/13/99 21:37 kuss 4He Iscan
  • 30924 12/13/99 16:50 Saha Arc-Energy measurement, Dec 11, 1999
  • 30923 12/13/99 14:08 R. Michaels XTrigMang ok again
  • 30922 12/13/99 13:28 R. Michaels reboot of hac
  • 30920 12/13/99 04:08 K. Aniol finish kapton
  • 30917 12/13/99 02:25 K. Aniol start run 1251
  • 30914 12/13/99 02:03 K. Aniol ep on kapton,845MeV
  • 30912 12/13/99 01:51 K. Aniol kin #17, Help!
  • 30911 12/13/99 00:07 Eric Voutier Shift summary
  • Sunday

  • 30881 12/12/99 17:43 R. Michaels calibration of SPOT++ checked
  • 30869 12/12/99 15:50 Calarco OTR
  • 30868 12/12/99 15:49 Calarco OTR
  • 30866 12/12/99 15:43 Calarco run plan, event rate
  • 30860 12/12/99 14:19 R. Michaels Notes re: usage of XTrigMang
  • 30859 12/12/99 13:21 Calarco magnet settings
  • 30858 12/12/99 13:18 quemener strip charts
  • 30857 12/12/99 13:17 quemener strip charts
  • 30856 12/12/99 12:56 R. Michaels new vs old XTrigMang
  • 30855 12/12/99 12:41 Calarco magnet settings
  • 30853 12/12/99 12:23 R. Michaels trigger problems, solutions
  • 30848 12/12/99 10:20 Calarco status
  • 30846 12/12/99 08:50 A. Deur eP energy measurement
  • 30844 12/12/99 08:14 M. Rvachev Optics 845MeV
  • 30787 12/12/99 00:13
  • Saturday

  • 30785 12/11/99 22:36 epstein strip chart
  • 30784 12/11/99 22:35 epstein strip chart
  • 30783 12/11/99 18:45 epstein strip chart
  • 30782 12/11/99 18:40 epstein strip chart
  • 30781 12/11/99 18:38 epstein strip chart
  • 30779 12/11/99 16:16 Kox/Quemener Shift Summary 12-11-99 / 8-16h
  • 30777 12/11/99 14:08 quemener strip charts
  • 30776 12/11/99 14:06 quemener strip charts
  • Friday

    
    

    Thursday

  • 30770 12/09/99 17:08 Jack Dead fan tray.
  • Wednesday

  • 30767 12/08/99 09:53 R. Michaels BPM Pedestal Run 1191
  • Tuesday

  • 30762 12/07/99 18:33 M Jones added deadtime to runend
  • Monday

  • 30760 12/06/99 16:49 R. Michaels ped run works again
  • 30756 12/06/99 13:25 R. Michaels pedestal runs problems
  • 30755 12/06/99 13:23 R. Michaels New ROC13 Data
  • 30727 12/06/99 09:58 M Jones Modified HAC .login
  • 30726 12/06/99 09:56 M Jones modified fastEpicsLogger
  • 30722 12/06/99 07:52 nilanga 845 optics
  • 30712 12/06/99 07:03 Eric Voutier Shift summary
  • 30686 12/06/99 04:39 Saha No Arc-energy measurement
  • 30685 12/06/99 02:08 saha Fast Feedback in Dispersive optics
  • Sunday

  • 30683 12/05/99 22:12 M. Rvachev Cycling field values
  • 30672 12/05/99 19:56 M. Rvachev Harp scan
  • 30670 12/05/99 19:52 M. Rvachev Harp scan
  • 30657 12/05/99 18:05 M. Rvachev Harp scan
  • 30652 12/05/99 17:49 quemener strip charts
  • 30643 12/05/99 16:58 quemener strip charts
  • 30640 12/05/99 16:26 M. Rvachev Harp scan
  • 30614 12/05/99 13:30 A. Deur eP energy measurement
  • 30613 12/05/99 12:59 kuss fast feedback in dispersive mode is fixed
  • 30610 12/05/99 06:22 M. Rvachev Harp scans
  • 30592 12/05/99 04:19 Saha Standard tunes for Hall A at 845 MeV
  • Saturday

  • 30585 12/04/99 23:34 M. Rvachev Quad cycling done
  • 30582 12/04/99 18:09 kuss fast feedback repaired
  • 30580 12/04/99 14:17 M Jones modified fastEpicsLogger
  • 30578 12/04/99 11:45 Saha Hall A Beam Checks for e89044
  • 30577 12/04/99 11:31 kuss John's quad cycle procedure
  • 30576 12/04/99 10:57 Saha Harp scans vs. "Spot"
  • 30575 12/04/99 10:03 R. Michaels spot++ needs check
  • 30574 12/04/99 09:16 M. Rvachev Shift summary
  • 30570 12/04/99 08:24 M.Rvachev Search for "inverted" optics
  • 30561 12/04/99 07:08 M. Rvachev Run comments messed up
  • 30557 12/04/99 06:58 M. Rvachev Magnets screen
  • 30554 12/04/99 06:47 M.Rvachev Momentum change to -2%
  • 30551 12/04/99 06:39 M. Rvachev BeO target
  • Friday

  • 30533 12/03/99 23:52 kuss energy striptools
  • 30534 12/03/99 23:52 kuss position stripchart
  • 30532 12/03/99 21:34 kuss beam parameter striptools
  • 30531 12/03/99 21:24 kuss access for resetting the magnets
  • 30530 12/03/99 15:25 M Jones modified Main epics display
  • Thursday

    
    

    Wednesday